Grade 11Algebra

Limx-0 ln(1+x²)/x . Solve by using both simple method and l hospital rule

Rohit
9 Years agoGrade 11

1 Answer

Tanishq Chourishi
9 Years ago
scine its 0/0 form directly apply the lohopitall theorem
 
(d(ln(1+x^2))/dx)/(dx/dx)={(1/(1+x^2))*(2x)}/1==>(2x)/(1+x^2)
 
now put lim x---->0
we get 2*0/(1+0)=0...
by general method 
lima--->0  {ln(1+a)}/a   =1
{ln{1+x2}/x2}*x2}/x={ln{1+x2}/x2}*x=1*x=1*0=0